# Text-Trained LLMs Can Zero-Shot Extrapolate PDE Dynamics

### v1

@@ A.11 Token-Level Distribution Visualizations Across Learning Stages
Figure 21: Representative token distributions across ICL stages, extracted from the Llama-3.1-8B model’s softmax outputs on a randomly sampled initial condition (same as in the multi-step rollout example). For clarity, only the top 8 candidate tokens (by probability) are shown per spatial position. (a) Syntax-only stage: separator tokens (e.g., commas) are predicted with near-perfect confidence, while spatial values are either deterministic but incorrect or act as generic placeholders. (b) Exploratory stage: spatial value distributions broaden, reflecting increased uncertainty and competing hypotheses with partial alignment to ground truth. (c) Consolidation stage: uncertainty decreases, and distributions sharpen around the true target values.

### v2

@@ 11 Token-Level Distribution Visualizations Across Learning Stages
Figure S20: Representative token distributions across ICL stages, extracted from the Llama-3.1-8B model’s softmax outputs on a randomly sampled initial condition (same as in the multi-step rollout example). For clarity, only the top 8 candidate tokens (by probability) are shown per spatial position. (a) Syntax-only stage: separator tokens (e.g., commas) are predicted with near-perfect confidence, while spatial values are either deterministic but incorrect or act as generic placeholders. (b) Exploratory stage: spatial value distributions broaden, reflecting increased uncertainty and competing hypotheses with partial alignment to ground truth. (c) Consolidation stage: uncertainty decreases, and distributions sharpen around the true target values.